### v1

@@ 6 Conclusion
In portfolio experiments, the proposed method enables global minimum-variance (GMV) strategies to achieve lower out-of-sample volatility with moderate turnover, underlining the tangible economic value of improved covariance estimation. By bridging cutting-edge DL techniques with established financial principles, our work highlights the promise of sophisticated spatiotemporal modelling for risk management and allocation decisions, especially at horizons where structural shifts and evolving correlations pose unique forecasting challenges.
Still, several research avenues warrant further exploration. First, extending the network architecture to accommodate skewed or heavy-tailed distributions could better capture tail risk in times of market stress. Second, explicitly incorporating transaction costs or liquidity constraints within the optimisation process might yield more realistic and implementable trading strategies. Finally, evaluating the model in larger cross-sectional settings or applied to alternative asset classes (e.g., commodities, crypto assets) would illuminate its scalability and robustness.
@@ A.2 Hyperparameter Grid Search
A.3 Extended Data Tables
@@
Table A4: Performance of the GMV portfolios constructed using the covariance forecasts from the 11 models under consideration plus the equal-weighted portfolio for the different market regimes. The portfolios are compared based on their annualised out-of-sample standard deviation ($\sigma_{m}$) and average out-of-sample turnover ($\text{TO}_{m}$), respectively.
Daily Weekly Monthly
$\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$ $\sigma_{m}^{2}$ $TO_{m}$
Bull-1 $1/N$ 0.009883 0.005169 0.009873 0.002335 0.009907 0.001103
$NA$ 0.000543 0.150112 0.000794 0.060372 0.000834 0.026885
$NA^{F}$ 0.000731 0.001378 0.000732 0.000607 0.000736 0.000327
$EWMA$ 0.000735 0.088541 0.000748 0.043884 0.000754 0.015613
$LW$ 0.001054 0.091979 0.001572 0.044883 0.001692 0.022283
$LW^{F}$ 0.000780 0.001228 0.000781 0.000610 0.000783 0.000349
$CCC$ 0.000620 0.119486 0.000681 0.053221 0.000728 0.013605
$DCC$ 0.000529 0.207262 0.000698 0.093558 0.000781 0.024324
$DCC^{NL}$ 0.000522 0.217432 0.000784 0.097809 0.000875 0.034673
$PCA$ 0.000850 0.236835 0.000827 0.090803 0.000795 0.020906
$RMT$ 0.000856 0.150601 0.000860 0.067402 0.000826 0.025859
$CAB$ 0.000634 0.086489 0.000656 0.055931 0.000707 0.023736
Bear $1/N$ 0.029939 0.007221 0.029940 0.003443 0.029793 0.001599
$NA$ 0.002207 0.092111 0.002682 0.032018 0.002959 0.018142
$NA^{F}$ 0.002891 0.001650 0.002892 0.000935 0.002906 0.000436
$EWMA$ 0.002908 0.079065 0.002843 0.036314 0.003120 0.019997
$LW$ 0.003118 0.095723 0.003970 0.041663 0.004754 0.021610
$LW^{F}$ 0.003196 0.002035 0.003197 0.001043 0.003201 0.000570
$CCC$ 0.002497 0.129487 0.002646 0.043843 0.002590 0.012451
$DCC$ 0.002355 0.067768 0.002639 0.015372 0.002636 0.005911
$DCC^{NL}$ 0.002269 0.083567 0.002652 0.024277 0.002730 0.010308
$PCA$ 0.002800 0.114256 0.002792 0.043637 0.002864 0.011111
$RMT$ 0.002983 0.094396 0.002896 0.037508 0.003153 0.023308
$CAB$ 0.002273 0.030506 0.002464 0.022686 0.002656 0.016836
Bull-2 $1/N$ 0.019967 0.005666 0.019956 0.002382 0.019933 0.001334
$NA$ 0.002823 0.118997 0.003526 0.053457 0.003667 0.029686
$NA^{F}$ 0.003398 0.001426 0.003400 0.000577 0.003406 0.000261
$EWMA$ 0.003432 0.073812 0.003433 0.034460 0.003630 0.016887
$LW$ 0.003512 0.102531 0.004274 0.048313 0.004245 0.020765
$LW^{F}$ 0.003874 0.002113 0.003879 0.001123 0.003889 0.000788
$CCC$ 0.003233 0.129874 0.003635 0.047812 0.003671 0.015276
$DCC$ 0.002947 0.071444 0.003294 0.026871 0.003364 0.009476
$DCC^{NL}$ 0.002924 0.085242 0.003318 0.034070 0.003365 0.012661
$PCA$ 0.003793 0.194898 0.003678 0.063899 0.003597 0.024379
$RMT$ 0.003668 0.118539 0.003720 0.055498 0.003762 0.029837
$CAB$ 0.003323 0.087821 0.003547 0.052596 0.003885 0.029487

### v2

@@ 6 Conclusion
In portfolio experiments, the proposed method enables GMV strategies to achieve lower out-of-sample volatility with moderate turnover, underlining the tangible economic value of improved covariance forecast. By bridging cutting-edge DL techniques with established financial principles, our work highlights the promise of sophisticated spatiotemporal modelling for risk management and allocation decisions, especially at horizons where structural shifts and evolving correlations pose unique forecasting challenges.
Still, two research avenues warrant further exploration. First, extending the network architecture to accommodate skewed or heavy-tailed distributions could better capture tail risk in times of market stress. Second, explicitly incorporating transaction costs or liquidity constraints within the optimisation process might yield more realistic and implementable trading strategies.
@@ A.2 Hyperparameter Grid Search
A.3 Extended ETF List
Table A2: List of ETFs used in the larger dataset.
Name Ticker Incept. Date Asset Class
iShares Core International Aggregate Bond ETF IAGG 10/11/2015 Fixed Income
iShares MSCI Saudi Arabia ETF KSA 16/09/2015 Equity
iShares Interest Rate Hedged Long-Term Corporate Bond ETF IGBH 22/07/2015 Fixed Income
iShares Currency Hedged MSCI ACWI ex U.S. ETF HAWX 29/06/2015 Equity
iShares Currency Hedged MSCI EAFE Small-Cap ETF HSCZ 29/06/2015 Equity
iShares MSCI Intl Value Factor ETF IVLU 16/06/2015 Equity
iShares Convertible Bond ETF ICVT 02/06/2015 Fixed Income
iShares U.S. Equity Factor ETF LRGF 28/04/2015 Equity
iShares U.S. Small-Cap Equity Factor ETF SMLF 28/04/2015 Equity
iShares International Equity Factor ETF INTF 28/04/2015 Equity
iShares International Small-Cap Equity Factor ETF ISCF 28/04/2015 Equity
iShares Global Equity Factor ETF GLOF 28/04/2015 Equity
iShares Exponential Technologies ETF XT 19/03/2015 Equity
iShares iBonds Dec 2025 Term Corporate ETF IBDQ 11/03/2015 Fixed Income
iShares Short Maturity Municipal Bond Active ETF MEAR 03/03/2015 Fixed Income
iShares U.S. Fixed Income Balanced Risk Systematic ETF FIBR 24/02/2015 Fixed Income
iShares MSCI Intl Quality Factor ETF IQLT 13/01/2015 Equity
iShares MSCI Intl Momentum Factor ETF IMTM 13/01/2015 Equity
iShares MSCI ACWI Low Carbon Target ETF CRBN 08/12/2014 Equity
iShares GSCI Commodity Dynamic Roll Strategy ETF COMT 15/10/2014 Commodity
iShares Currency Hedged MSCI Emerging Markets ETF HEEM 23/09/2014 Equity
iShares Currency Hedged MSCI Eurozone ETF HEZU 09/07/2014 Equity
iShares Global REIT ETF REET 08/07/2014 Real Estate
iShares Core Total USD Bond Market ETF IUSB 10/06/2014 Fixed Income
iShares Core Dividend Growth ETF DGRO 10/06/2014 Equity
iShares Core MSCI Europe ETF IEUR 10/06/2014 Equity
iShares Core MSCI Pacific ETF IPAC 10/06/2014 Equity
iShares Interest Rate Hedged Corporate Bond ETF LQDH 27/05/2014 Fixed Income
iShares Interest Rate Hedged High Yield Bond ETF HYGH 27/05/2014 Fixed Income
iShares MSCI UAE ETF UAE 29/04/2014 Equity
iShares MSCI Qatar ETF QAT 29/04/2014 Equity
iShares Yield Optimized Bond ETF BYLD 22/04/2014 Fixed Income
iShares Treasury Floating Rate Bond ETF TFLO 03/02/2014 Fixed Income
iShares Currency Hedged MSCI EAFE ETF HEFA 31/01/2014 Equity
iShares Currency Hedged MSCI Japan ETF HEWJ 31/01/2014 Equity
iShares Ultra Short-Term Bond Active ETF ICSH 11/12/2013 Fixed Income
iShares 0-5 Year High Yield Corporate Bond ETF SHYG 15/10/2013 Fixed Income
iShares 0-5 Year Investment Grade Corporate Bond ETF SLQD 15/10/2013 Fixed Income
iShares Short Duration Bond Active ETF NEAR 25/09/2013 Fixed Income
iShares MSCI USA Quality Factor ETF QUAL 16/07/2013 Equity
iShares MSCI USA Momentum Factor ETF MTUM 16/04/2013 Equity
iShares MSCI USA Value Factor ETF VLUE 16/04/2013 Equity
iShares MSCI USA Size Factor ETF SIZE 16/04/2013 Equity
iShares Core MSCI EAFE ETF IEFA 18/10/2012 Equity
iShares Core MSCI Emerging Markets ETF IEMG 18/10/2012 Equity
iShares Core MSCI Total International Stock ETF IXUS 18/10/2012 Equity
iShares Core 1-5 Year USD Bond ETF ISTB 18/10/2012 Fixed Income
iShares J.P. Morgan EM Corporate Bond ETF CEMB 17/04/2012 Fixed Income
iShares J.P. Morgan EM High Yield Bond ETF EMHY 03/04/2012 Fixed Income
iShares US & Intl High Yield Corp Bond ETF GHYG 03/04/2012 Fixed Income
iShares Morningstar Multi-Asset Income ETF IYLD 03/04/2012 Multi Asset
iShares International High Yield Bond ETF HYXU 03/04/2012 Fixed Income
iShares Emerging Markets Dividend ETF DVYE 23/02/2012 Equity
iShares Asia/Pacific Dividend ETF DVYA 23/02/2012 Equity
iShares U.S. Treasury Bond ETF GOVT 14/02/2012 Fixed Income
iShares Aaa - A Rated Corporate Bond ETF QLTA 14/02/2012 Fixed Income
iShares CMBS ETF CMBS 14/02/2012 Others
iShares GNMA Bond ETF GNMA 14/02/2012 Fixed Income
iShares MSCI India Small-Cap ETF SMIN 08/02/2012 Equity
iShares MSCI Emerging Markets Asia ETF EEMA 08/02/2012 Equity
iShares MSCI India ETF INDA 02/02/2012 Equity
iShares MSCI Global Gold Miners ETF RING 31/01/2012 Equity
iShares MSCI Global Metals & Mining Producers ETF PICK 31/01/2012 Equity
iShares MSCI Global Silver and Metals Miners ETF SLVP 31/01/2012 Equity
iShares MSCI Agriculture Producers ETF VEGI 31/01/2012 Equity
iShares MSCI Global Energy Producers ETF FILL 31/01/2012 Equity
iShares MSCI Denmark ETF EDEN 25/01/2012 Equity
iShares MSCI United Kingdom Small-Cap ETF EWUS 25/01/2012 Equity
iShares MSCI Finland ETF EFNL 25/01/2012 Equity
iShares MSCI Norway ETF ENOR 23/01/2012 Equity
iShares MSCI World ETF URTH 10/01/2012 Equity
iShares MSCI USA Min Vol Factor ETF USMV 18/10/2011 Equity
iShares MSCI EAFE Min Vol Factor ETF EFAV 18/10/2011 Equity
iShares MSCI Global Min Vol Factor ETF ACWV 18/10/2011 Equity
iShares MSCI Emerging Markets Min Vol Factor ETF EEMV 18/10/2011 Equity
iShares J.P. Morgan EM Local Currency Bond ETF LEMB 18/10/2011 Fixed Income
iShares MSCI Emerging Markets Small-Cap ETF EEMS 16/08/2011 Equity
iShares Floating Rate Bond ETF FLOT 14/06/2011 Fixed Income
iShares Core High Dividend ETF HDV 29/03/2011 Equity
iShares MSCI China ETF MCHI 29/03/2011 Equity
iShares 0-5 Year TIPS Bond ETF STIP 01/12/2010 Fixed Income
iShares MSCI Brazil Small-Cap ETF EWZS 28/09/2010 Equity
iShares MSCI Philippines ETF EPHE 28/09/2010 Equity
iShares MSCI China Small-Cap ETF ECNS 28/09/2010 Equity
iShares MSCI New Zealand ETF ENZL 01/09/2010 Equity
iShares MSCI Poland ETF EPOL 25/05/2010 Equity
iShares MSCI USA Equal Weighted ETF EUSA 05/05/2010 Equity
iShares MSCI Indonesia ETF EIDO 05/05/2010 Equity
iShares MSCI Ireland ETF EIRL 05/05/2010 Equity
iShares MSCI Europe Financials ETF EUFN 20/01/2010 Equity
iShares 10+ Year Investment Grade Corporate Bond ETF IGLB 08/12/2009 Fixed Income
iShares Core 10+ Year USD Bond ETF ILTB 08/12/2009 Fixed Income
iShares India 50 ETF INDY 18/11/2009 Equity
iShares Russell Top 200 Growth ETF IWY 22/09/2009 Equity
iShares Russell Top 200 Value ETF IWX 22/09/2009 Equity
iShares Russell Top 200 ETF IWL 22/09/2009 Equity
iShares MSCI Peru and Global Exposure ETF EPU 19/06/2009 Equity
iShares Emerging Markets Infrastructure ETF EMIF 16/06/2009 Equity
iShares International Treasury Bond ETF IGOV 21/01/2009 Fixed Income
iShares 1-3 Year International Treasury Bond ETF ISHG 21/01/2009 Fixed Income
iShares Short-Term National Muni Bond ETF SUB 05/11/2008 Fixed Income
iShares Agency Bond ETF AGZ 05/11/2008 Fixed Income
iShares Core Growth Allocation ETF AOR 04/11/2008 Multi Asset
iShares Core Aggressive Allocation ETF AOA 04/11/2008 Multi Asset
iShares Core Moderate Allocation ETF AOM 04/11/2008 Multi Asset
iShares Core Conservative Allocation ETF AOK 04/11/2008 Multi Asset
iShares MSCI All Country Asia ex Japan ETF AAXJ 13/08/2008 Equity
iShares Global Clean Energy ETF ICLN 24/06/2008 Equity
iShares Global Timber & Forestry ETF WOOD 24/06/2008 Equity
iShares MSCI ACWI ETF ACWI 26/03/2008 Equity
iShares MSCI ACWI ex U.S. ETF ACWX 26/03/2008 Equity
iShares MSCI Israel ETF EIS 26/03/2008 Equity
iShares MSCI Thailand ETF THD 26/03/2008 Equity
iShares MSCI Turkey ETF TUR 26/03/2008 Equity
iShares MSCI Japan Small-Cap ETF SCJ 20/12/2007 Equity
iShares J.P. Morgan USD Emerging Markets Bond ETF EMB 17/12/2007 Fixed Income
iShares MSCI EAFE Small-Cap ETF SCZ 10/12/2007 Equity
iShares Global Infrastructure ETF IGF 10/12/2007 Equity
iShares MSCI Kokusai ETF TOK 10/12/2007 Equity
iShares Asia 50 ETF AIA 13/11/2007 Equity
iShares MSCI Chile ETF ECH 12/11/2007 Equity
iShares International Developed Real Estate ETF IFGL 12/11/2007 Real Estate
iShares MSCI Europe Small-Cap ETF IEUS 12/11/2007 Equity
iShares MSCI BIC ETF BKF 12/11/2007 Equity
iShares California Muni Bond ETF CMF 04/10/2007 Fixed Income
iShares New York Muni Bond ETF NYF 04/10/2007 Fixed Income
iShares National Muni Bond ETF MUB 07/09/2007 Fixed Income
iShares International Select Dividend ETF IDV 11/06/2007 Equity
iShares Core U.S. REIT ETF USRT 01/05/2007 Real Estate
iShares Residential and Multisector Real Estate ETF REZ 01/05/2007 Real Estate
iShares Mortgage Real Estate ETF REM 01/05/2007 Real Estate
iShares iBoxx High Yield Corporate Bond ETF HYG 04/04/2007 Fixed Income
iShares Preferred and Income Securities ETF PFF 26/03/2007 Equity
iShares MBS ETF MBB 13/03/2007 Fixed Income
iShares 1-5 Year Investment Grade Corporate Bond ETF IGSB 05/01/2007 Fixed Income
iShares Short Treasury Bond ETF SHV 05/01/2007 Fixed Income
iShares 3-7 Year Treasury Bond ETF IEI 05/01/2007 Fixed Income
iShares 5-10 Year Investment Grade Corporate Bond ETF IGIB 05/01/2007 Fixed Income
iShares Broad USD Investment Grade Corporate Bond ETF USIG 05/01/2007 Fixed Income
iShares 10-20 Year Treasury Bond ETF TLH 05/01/2007 Fixed Income
iShares Intermediate Government/Credit Bond ETF GVI 05/01/2007 Fixed Income
iShares Government/Credit Bond ETF GBF 05/01/2007 Fixed Income
iShares MSCI KLD 400 Social ETF DSI 14/11/2006 Equity
iShares Global Industrials ETF EXI 12/09/2006 Equity
iShares Global Consumer Staples ETF KXI 12/09/2006 Equity
iShares Global Consumer Discretionary ETF RXI 12/09/2006 Equity
iShares Global Materials ETF MXI 12/09/2006 Equity
iShares Global Utilities ETF JXI 12/09/2006 Equity
iShares S&P GSCI Commodity-Indexed Trust GSG 10/07/2006 Commodity
iShares U.S. Aerospace & Defense ETF ITA 01/05/2006 Equity
iShares U.S. Medical Devices ETF IHI 01/05/2006 Equity
iShares U.S. Home Construction ETF ITB 01/05/2006 Equity
iShares U.S. Broker-Dealers & Securities Exchanges ETF IAI 01/05/2006 Equity
iShares U.S. Regional Banks ETF IAT 01/05/2006 Equity
iShares U.S. Insurance ETF IAK 01/05/2006 Equity
iShares U.S. Healthcare Providers ETF IHF 01/05/2006 Equity
iShares U.S. Pharmaceuticals ETF IHE 01/05/2006 Equity
iShares U.S. Oil & Gas Exploration & Production ETF IEO 01/05/2006 Equity
iShares U.S. Oil Equipment & Services ETF IEZ 01/05/2006 Equity
iShares Silver Trust SLV 21/04/2006 Commodity
iShares Micro-Cap ETF IWC 12/08/2005 Equity
iShares MSCI EAFE Value ETF EFV 01/08/2005 Equity
iShares MSCI EAFE Growth ETF EFG 01/08/2005 Equity
iShares MSCI USA ESG Select ETF SUSA 24/01/2005 Equity
iShares Gold Trust IAU 21/01/2005 Commodity
iShares China Large-Cap ETF FXI 05/10/2004 Equity
iShares Morningstar Mid-Cap Growth ETF IMCG 28/06/2004 Equity
iShares Morningstar Growth ETF ILCG 28/06/2004 Equity
iShares Morningstar U.S. Equity ETF ILCB 28/06/2004 Equity
iShares Morningstar Mid-Cap ETF IMCB 28/06/2004 Equity
iShares Morningstar Value ETF ILCV 28/06/2004 Equity
iShares Morningstar Mid-Cap Value ETF IMCV 28/06/2004 Equity
iShares Morningstar Small-Cap Growth ETF ISCG 28/06/2004 Equity
iShares Morningstar Small-Cap Value ETF ISCV 28/06/2004 Equity
iShares Morningstar Small-Cap ETF ISCB 28/06/2004 Equity
iShares Core S&P Total U.S. Stock Market ETF ITOT 20/01/2004 Equity
iShares TIPS Bond ETF TIP 04/12/2003 Fixed Income
iShares Select Dividend ETF DVY 03/11/2003 Equity
iShares U.S. Transportation ETF IYT 06/10/2003 Equity
iShares Core U.S. Aggregate Bond ETF AGG 22/09/2003 Fixed Income
iShares MSCI Emerging Markets ETF EEM 07/04/2003 Equity
iShares MSCI South Africa ETF EZA 03/02/2003 Equity
iShares 20+ Year Treasury Bond ETF TLT 22/07/2002 Fixed Income
iShares 7-10 Year Treasury Bond ETF IEF 22/07/2002 Fixed Income
iShares iBoxx Investment Grade Corporate Bond ETF LQD 22/07/2002 Fixed Income
iShares 1-3 Year Treasury Bond ETF SHY 22/07/2002 Fixed Income
iShares Global Healthcare ETF IXJ 13/11/2001 Equity
iShares Global Tech ETF IXN 12/11/2001 Equity
iShares Global Energy ETF IXC 12/11/2001 Equity
iShares Global Financials ETF IXG 12/11/2001 Equity
iShares Global Comm Services ETF IXP 12/11/2001 Equity
iShares MSCI Pacific ex Japan ETF EPP 25/10/2001 Equity
iShares Latin America 40 ETF ILF 25/10/2001 Equity
iShares JPX-Nikkei 400 ETF JPXN 23/10/2001 Equity
iShares North American Natural Resources ETF IGE 22/10/2001 Equity
iShares MSCI EAFE ETF EFA 14/08/2001 Equity
iShares Russell Mid-Cap ETF IWR 17/07/2001 Equity
iShares Russell Mid-Cap Growth ETF IWP 17/07/2001 Equity
iShares Russell Mid-Cap Value ETF IWS 17/07/2001 Equity
iShares Semiconductor ETF SOXX 10/07/2001 Equity
iShares Expanded Tech-Software Sector ETF IGV 10/07/2001 Equity
iShares U.S. Digital Infrastructure and Real Estate ETF IDGT 10/07/2001 Equity
iShares Expanded Tech Sector ETF IGM 13/03/2001 Equity
iShares Biotechnology ETF IBB 05/02/2001 Equity
iShares Cohen & Steers REIT ETF ICF 29/01/2001 Real Estate
iShares Global 100 ETF IOO 05/12/2000 Equity
iShares S&P 100 ETF OEF 23/10/2000 Equity
iShares MSCI Eurozone ETF EZU 25/07/2000 Equity
iShares Europe ETF IEV 25/07/2000 Equity
iShares Core S&P U.S. Growth ETF IUSG 24/07/2000 Equity
iShares Core S&P U.S. Value ETF IUSV 24/07/2000 Equity
iShares Russell 2000 Growth ETF IWO 24/07/2000 Equity
iShares Russell 2000 Value ETF IWN 24/07/2000 Equity
iShares S&P Mid-Cap 400 Growth ETF IJK 24/07/2000 Equity
iShares S&P Mid-Cap 400 Value ETF IJJ 24/07/2000 Equity
iShares S&P Small-Cap 600 Value ETF IJS 24/07/2000 Equity
iShares S&P Small-Cap 600 Growth ETF IJT 24/07/2000 Equity
iShares MSCI Brazil ETF EWZ 10/07/2000 Equity
iShares MSCI Taiwan ETF EWT 20/06/2000 Equity
iShares U.S. Real Estate ETF IYR 12/06/2000 Real Estate
iShares U.S. Healthcare ETF IYH 12/06/2000 Equity
iShares Dow Jones U.S. ETF IYY 12/06/2000 Equity
iShares U.S. Consumer Discretionary ETF IYC 12/06/2000 Equity
iShares U.S. Financial Services ETF IYG 12/06/2000 Equity
iShares U.S. Industrials ETF IYJ 12/06/2000 Equity
iShares U.S. Energy ETF IYE 12/06/2000 Equity
iShares U.S. Consumer Staples ETF IYK 12/06/2000 Equity
iShares U.S. Utilities ETF IDU 12/06/2000 Equity
iShares U.S. Basic Materials ETF IYM 12/06/2000 Equity
iShares Russell 1000 Growth ETF IWF 22/05/2000 Equity
iShares Core S&P Mid-Cap ETF IJH 22/05/2000 Equity
iShares Core S&P Small-Cap ETF IJR 22/05/2000 Equity
iShares Russell 2000 ETF IWM 22/05/2000 Equity
iShares Russell 1000 Value ETF IWD 22/05/2000 Equity
iShares S&P 500 Growth ETF IVW 22/05/2000 Equity
iShares S&P 500 Value ETF IVE 22/05/2000 Equity
iShares Russell 3000 ETF IWV 22/05/2000 Equity
iShares U.S. Financials ETF IYF 22/05/2000 Equity
iShares U.S. Telecommunications ETF IYZ 22/05/2000 Equity
iShares Core S&P 500 ETF IVV 15/05/2000 Equity
iShares Russell 1000 ETF IWB 15/05/2000 Equity
iShares U.S. Technology ETF IYW 15/05/2000 Equity
iShares MSCI South Korea ETF EWY 09/05/2000 Equity
iShares MSCI Japan ETF EWJ 12/03/1996 Equity
iShares MSCI United Kingdom ETF EWU 12/03/1996 Equity
iShares MSCI Canada ETF EWC 12/03/1996 Equity
iShares MSCI Australia ETF EWA 12/03/1996 Equity
iShares MSCI Mexico ETF EWW 12/03/1996 Equity
iShares MSCI Switzerland ETF EWL 12/03/1996 Equity
iShares MSCI Germany ETF EWG 12/03/1996 Equity
iShares MSCI Spain ETF EWP 12/03/1996 Equity
iShares MSCI Singapore ETF EWS 12/03/1996 Equity
iShares MSCI Hong Kong ETF EWH 12/03/1996 Equity
iShares MSCI France ETF EWQ 12/03/1996 Equity
iShares MSCI Italy ETF EWI 12/03/1996 Equity
iShares MSCI Sweden ETF EWD 12/03/1996 Equity
iShares MSCI Malaysia ETF EWM 12/03/1996 Equity
iShares MSCI Netherlands ETF EWN 12/03/1996 Equity
iShares MSCI Austria ETF EWO 12/03/1996 Equity
iShares MSCI Belgium ETF EWK 12/03/1996 Equity
A.4 Extended Tables